Rectangle RECT with the coordinates as follows:
R (-2, 1) (bottom left corner of rectangle)
E (-2, 6) (top left corner of rectangle)
C (8,6) (top right corner of rectangle)
T (8,1) (bottom right corner of rectangle)
Find the lengths of the sides . Then find the area.

Answer by checkley77(12569)

The x measurement=8+2=10
The y measurement=6-1=5
Area=10*5=50